OpenStudy (anonymous):

simplify expression 2/(2+[radical]3)

OpenStudy (anonymous):

\[\frac{ 2 }{ 2+\sqrt{3} }\]

hartnn (hartnn):

multiply and divide by 2-\(\sqrt 3\)

hartnn (hartnn):

what u get ?

OpenStudy (anonymous):

I got 2[radical]3 + 4 ): but it's not an option..

hartnn (hartnn):

because thats not correct.

hartnn (hartnn):

\(\huge \frac{2(2-\sqrt3)}{(2+\sqrt3)(2-\sqrt3)}=?\) what does denominator simplify to ?

hartnn (hartnn):

use \((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

so the denominator is one.

hartnn (hartnn):

yes.

hartnn (hartnn):

what about numerator ?

OpenStudy (anonymous):

why is it \[4-2\sqrt{3}\]?

hartnn (hartnn):

yes, absolutely correct. isn't that an option ?

OpenStudy (anonymous):

yes, thank you!

hartnn (hartnn):

welcome ^_^